# Boosting Semi-Supervised 2D Human Pose Estimation by Revisiting Data Augmentation and Consistency Training

### v2

@@ 5.4 Ablation Studies
Three advanced augmentations are selected: RandAugment [11], TrivialAugment [38] and YOCO [17]. We refer to them as $T_{RA}$, $T_{TA}$ and $T_{YOCO}$. For $T_{YOCO}$, it may be based on $T_{RA}$ or $T_{TA}$. And we compare them with previous SOTA augmentations $T_{JO}$ and $T_{JC}$ for SSHPE, and our recommended $T_{JOCO}$ and $T_{JCCM}$. We did not compare to Mixup families [70, 21, 43, 36] or AutoAug families [10, 33, 18, 72]. Because Mixup is verified not to work with SSHPE, and AutoAug needs to search optimal parameters. Finally, as shown in Tab. 5.4 and Fig. 5.4, our optimal combinations are always better than $T_{RA}$-based or $T_{TA}$-based $T_{YOCO}$, which are meticulously designed and also composed of existing basic augmentations. These further prove the superiority and conciseness of our synergistic combinations.
5.4.2 Training Techniques of Multiple Heatmaps.
In this part, we present additional techniques commonly used for unsupervised consistency training. Especially, for predicted multi-heatmaps, we propose to optimize them by applying the multi-loss learning (ML) as in Eq. 4. Other two alternative techniques are confidence masking (CM) and heatmaps fusion (HF). For CM, the consistency loss term in each mini-batch is computed only on keypoint channels whose maximum activation value is greater than a threshold $\tau$, which is set as 0.5. For HF, also termed as heatmaps ensemble, we sum and average multi-path heatmaps to obtain a fused heatmap for loss computing. Then, we compare MultiAugs (e.g., $m_{11}$ and $m_{12}$) using either of these three techniques ML, CM and HF.
As shown in Fig. 11, our ML is strictly superior than the other two techniques under either $m_{11}$ or $m_{12}$. For CM, we assume it may filter out some keypoint heatmaps with low confidence but high quality. This surely has a negative impact. For HF, although it is widely used in other SSL tasks for model ensemble, it may not necessarily be applicable to our intermediate keypoint heatmaps. We deem this is because each predicted heatmap is distinctive and meaningful (see Fig. 11). It is tricky to replace them equivalently with a fused heatmap. In comparison, our multi-loss learning is a simple yet effective choice.
@@
Figure 10: The convergence curves and best mAP results of two MultiAugs schemes with using three different training techniques.
Figure 10: The convergence curves and best mAP results of two MultiAugs schemes with using three different training techniques.
@@
6 Conclusions
@@
In this paper, we aim to boost semi-supervised human pose estimation (SSHPE) from two perspectives: data augmentation and consistency training. Instead of inventing advanced augmentations in isolation, we attempt to synergistically utilize existing augmentations, and handily generate superior ones by novel combination paradigms. The discovered collaborative combinations have intuitive interpretability. We verified their advantages in solving the SSHPE problem. For consistency training, we abandon the convention of stacking networks to increase unsupervised losses, and train a single network by optimizing multi-path consistency losses for the same batch of unlabeled images. Combined with the optimal hard augmentations set, this plain and compact strategy is proven to be effective, and leads to better performance on public benchmarks.
Last but not least, we declare that the synergy effects of augmentations and multiple consistency losses are generic and generalizable for other SSL vision tasks such as image classification, object detection and semantic segmentation. We will explore them in the future.

### v3

@@ V-C1 Comparing to Other Specially Designed Advanced Augmentations
Three advanced augmentations are selected: RandAugment [26], TrivialAugment [31] and YOCO [34]. We refer to them as $T_{RA}$, $T_{TA}$ and $T_{YOCO}$. For $T_{YOCO}$, it may be based on $T_{RA}$ or $T_{TA}$. And we compare them with previous SOTA augmentations $T_{JO}$ and $T_{JC}$ for SSHPE, and our recommended $T_{JOCO}$ and $T_{JCCM}$. We did not compare to Mixup families [23, 30, 32, 33] or AutoAug families [25, 27, 28, 29]. Because Mixup is verified not to work with SSHPE, and AutoAug needs to search optimal parameters. Finally, as shown in Tab. V-C and Fig. V-C, our optimal combinations are always better than $T_{RA}$-based or $T_{TA}$-based $T_{YOCO}$, which are meticulously designed and also composed of existing basic augmentations. These further prove the superiority and conciseness of our synergistic combinations.
@@
Figure 10: The convergence curves and best mAP results of two MultiAugs schemes when using different training techniques.
@@
V-C2 Training Techniques of Multiple Heatmaps
In this part, we present additional techniques commonly used for unsupervised consistency training. Especially, for predicted multi-heatmaps, we propose to optimize them by applying the multi-loss learning (ML) as in Eq. 4. Other two alternative techniques are confidence masking (CM) and heatmaps fusion (HF). For CM, the consistency loss term in each mini-batch is computed only on keypoint channels whose maximum activation value is greater than a threshold $\tau$, which is set as 0.5. For HF, also termed as heatmaps ensemble, we sum and average multi-path heatmaps to obtain a fused heatmap for loss computing. Then, we compare MultiAugs (e.g., $m_{11}$ and $m_{12}$) using either of these three techniques ML, CM and HF. As shown in Fig. 10, our ML is strictly superior than the other two techniques under either $m_{11}$ or $m_{12}$. For CM, we assume it may filter out some keypoint heatmaps with low confidence but high quality. This surely has a negative impact. For HF, although it is widely used in other SSL tasks for model ensemble, it may not necessarily be applicable to our intermediate keypoint heatmaps. We deem this is because each predicted heatmap is distinctive and meaningful (please refer Fig. 11 for details). It is tricky to replace them equivalently with a fused heatmap. In comparison, our multi-loss learning is a simple yet effective choice.
TABLE XVI: AP results of baseline methods after increasing the batch size. The used backbone is ResNet18.
  Method Nets Loss Num. Batch Size Labeled Samples
1K 5K 10K
PoseCons 1 1 32 42.1 52.3 57.3
PoseCons 1 1 128 42.3 52.6 57.5
PoseDual 2 1 32 44.6 55.6 59.6
PoseDual 2 1 128 44.9 58.7 59.6
TABLE XVII: AP results of our methods after adjusting the augmentation way of inputs. The used backbone is ResNet18.
  Method Nets Input Style Labeled Samples
1K 5K 10K
Ours (Single) 1 1-vs-n 45.5 56.2 59.9
Ours (Single) 1 n-vs-n 45.6 56.4 59.8
Ours (Dual) 2 1-vs-n 49.7 58.8 61.8
Ours (Dual) 2 n-vs-n 49.7 58.9 61.9
TABLE XVIII: Quantitative comparison of training time. The integer with the marker # means multi-path losses number.
  Method PoseCons Ours (Single,2#) Ours (Single,3#) Ours (Single,4#)
Time $T_{0}$ 1.36*$T_{0}$ 1.50*$T_{0}$ 1.83*$T_{0}$
Method PoseDual Ours (Dual,2#) Ours (Dual,3#) Ours (Dual,4#)
Time 2.49*$T_{0}$ 2.62*$T_{0}$ 2.88*$T_{0}$ 3.14*$T_{0}$
V-C3 Influence of Batch Size
We here investigate whether the multi-path consistency loss is sensitive to the training batch size. In fact, when performing ablation studies in Tab. III-A3 using the single-path loss and Tab. III-B2 using multi-path losses, we always set a fixed batch size 32. Moreover, for comparative experiments in Tab. V, we still keep the batch size as 32 and use the optimal four-path losses. Now, to investigate the possible impact of different batch sizes, we report the effects of PoseCons and PoseDual when the batch size is 128. As can be seen in Tab. XVI, after increasing the batch size of PoseCons or PoseDual accordingly, the final mAP results under different labeling rates (e.g., 1K, 5K and 10K) did not get significantly better. This indicates that batch size does not have a large impact on the performance of existing methods.
V-C4 Influence of Easy Augmentations
We also need to probe whether to use a single constant easy augmentation as input for multi-path losses (the pair $\{I_{e}\}$ + $\{I_{h_{1}},...,I_{h_{n}}\}$, termed as 1-vs-n) or to use different easy augmentations multiple times as input (the pair $\{I_{e_{1}}$,…,$I_{e_{n}}\}$ + {$I_{h_{1}}$,…,$I_{h_{n}}\}$, termed as n-vs-n). As shown in Tab. XVII, whether using 1-v-n augmented input or n-vs-n augmented input, the final mAP results obtained under various labeling rates are not significantly different. This is mainly because the used easy augmentation is always fixed (e.g., $T_{A30}$), so the input does not change in essence when applying 1-vs-n input or n-vs-n. These also reveal the necessity and rationality for us to focus on discussing how to find more advanced hard augmentations.
V-C5 Training Efficiency of Our Method
Finally, in order to fairly and reasonably reflect the efficiency of our method in the training phase, we compare the training time of our method with that of PoseCons and PoseDual. The strong augmentation used by them is $T_{JC}$. Considering that our method often uses different strong augmentations, their computation is not the main bottleneck. Therefore, all strong augmentations in our method are also replaced into $T_{JC}$. Assuming that the total training time of PoseCons is one unit time $T_{0}$, which is actually about 7 hours. Then the total training time of running other methods is summarized in Tab. XVIII. We can see that when using four-path losses, although the training time increases, it is still faster than PoseDual (1.83*$T_{0}$ vs. 2.49*$T_{0}$). Referring to the quantitative results in Tab. V, our method using a single-network and four-path losses achieves higher mAP than PoseDual. In addition, when using a dual-network with four-path losses, the total training time does not increase significantly (2.49*$T_{0}$ vs. 3.14*$T_{0}$). These indicate that our method is both efficient and effective.
VI Conclusions
@@
In this paper, we aim to boost semi-supervised human pose estimation (SSHPE) from two perspectives: data augmentation and consistency training. Instead of inventing advanced augmentations in isolation, we attempt to synergistically utilize existing augmentations, and handily generate superior ones by novel combination paradigms. The discovered collaborative combinations have intuitive interpretability. We verified their advantages in solving the SSHPE problem. For consistency training, we abandon the convention of stacking networks to increase unsupervised losses, and train a single network by optimizing multi-path consistency losses for the same batch of unlabeled images. Combined with the optimal hard augmentations set, this plain and compact strategy is proven to be effective, and leads to better performance on various public benchmarks across different human keypoint domains. Last but not least, we declare that the synergy effects of augmentations and multiple consistency losses are generic and generalizable for other SSL vision tasks such as image classification, object detection, semantic segmentation and 3D regression. We will explore them in the future.